### v1

# Privacy-Preserving Dataset Combination

Abstract.
Access to diverse, high-quality datasets is crucial for machine learning model performance, yet data sharing remains limited by privacy concerns and competitive interests, particularly in regulated domains like healthcare. This dynamic especially disadvantages smaller organizations that lack resources to purchase data or negotiate favorable sharing agreements. We present SecureKL, a privacy-preserving framework that enables organizations to identify beneficial data partnerships without exposing sensitive information. Building on recent advances in dataset combination methods, we develop a secure multiparty computation protocol that maintains strong privacy guarantees while achieving ¿90% correlation with plaintext evaluations. In experiments with real-world hospital data, SecureKL successfully identifies beneficial data partnerships that improve model performance for intensive care unit mortality prediction while preserving data privacy. Our framework provides a practical solution for organizations seeking to leverage collective data resources while maintaining privacy and competitive advantages. These results demonstrate the potential for privacy-preserving data collaboration to advance machine learning applications in high-stakes domains while promoting more equitable access to data resources. Our code is publicly available at https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Empirical scaling laws have established clear relationships between model performance and three key factors: compute, data, and model size (Hestness2017DeepLS, ; Kaplan2020ScalingLF, ). These relationships have driven remarkable improvements across computer vision, language processing, speech recognition, reinforcement learning, and healthcare applications (sun2017revisiting, ; mahajan2018exploring, ; brown2020language, ; hoffmann2022training, ; park2019specaugment, ; zhang2020pushing, ; ashvin2020accelerating, ; reed2022generalist, ; sheller2020federated, ; mckinney2020international, ). Beyond simple scaling, the diversity of training data has proven crucial for enhancing model robustness to distribution shifts and mitigating performance disparities across demographic groups (shen2024data, ; miller2021accuracy, ).
However, access to data varies significantly across entities and domains. While large tech companies have the data and compute resources to train the foundational models that now dominate general tasks, smaller players often lack such access. Domain-specific data is also becoming increasingly valuable for fine-tuning these general models (alsentzer2019publicly, ; gururangan2020don, ; lee2020biobert, ), creating a competitive advantage for the data owners. As a result, entities with domain-specific data are more reluctant to share it for free, opting instead to sell it in emerging data markets (acemoglu2022too, ; huang2021toward, ; liang2018survey, ). This dynamic particularly disadvantages smaller organizations, which often lack both the resources to purchase data and the leverage to negotiate favorable sharing agreements.
Scaling datasets in regulated domains is particularly challenging due to legal constraints and unpredictable outcomes from altering the training data composition. In healthcare, for example, patient data is heavily regulated by laws such as the Health Insurance Portability and Accountability Act (HIPAA), which imposes strict data-sharing constraints to protect patient privacy. Moreover, accumulating data from multiple sources introduces the risk of domain shift, where data from different distributions may degrade model performance instead of improving it (koh2021wilds, ; wang2022generalizing, ; taori2020measuring, ). Contrary to the intuition that more data always leads to better performance, combining datasets does not guarantee improvements — in fact, performance can decrease when incorporating additional data sources (bradley2021unrepresentative, ; meng2018statistical, ; shen2024data, ). This non-monotonic behavior means that carefully selecting which datasets to combine is crucial, as using all available data may actually perform worse than using an optimal subset.
While existing methods aim to identify datasets combinations that can improve performance, they often assume access to all relevant datasets (shen2024data, ; ilyas2022datamodels, ). However, this is impractical because entities are reluctant to share their data due to privacy risks or the competitive value they associate with it. This reluctance to share creates a bottleneck in improving model performance. In high-stakes settings like healthcare, this bottleneck is particularly detrimental, as access to diverse data has the potential to drive significant advancements in patient care and outcomes (beam2018big, ; ghassemi2019practical, ; chen2020treating, ; beaulieu2019trends, ).
To address these challenges, we propose SecureKL, a privacy-preserving approach for guiding dataset combinations without requiring direct data or model sharing. Our work makes three primary technical contributions:
  1. (1)
    We introduce a framework for practical secure data combination that enables organizations to evaluate potential partnerships while maintaining data privacy. Our framework categorizes existing approaches by their privacy leakage risks and provides a systematic way to assess the trade-offs between data utility and privacy preservation.
  2. (2)
    We extend KL-XY score (shen2024data, ) with a secure multiparty protocol that enables privacy-preserving evaluation of potential data partnerships. Our protocol maintains strong privacy guarantees while utilizing the complete underlying datasets, achieving ¿90% correlation with plaintext computations.
  3. (3)
    Through extensive evaluation in low-data, high-stakes settings, we demonstrate that our method successfully identifies beneficial data partnerships for intensive care unit (ICU) mortality prediction, improving classifier performance for the source hospital.
  4. (4)
    In simulated scenarios requiring selection of three partner hospitals, Private-KL-XY outperforms alternative selection strategies including demographic-based selection (using gender, race, and age), and limited-sample plaintext selection.
We argue that our method presents an appealing trade-off of privacy and utility by preserving privacy for both parties while using all the underlying data.
2. Problem Setup
Consider a binary prediction task for ICU patient mortality based on electronic medical records. A source hospital $H_{o}$ has historical patient data $\mathcal{D}_{o}$ containing static past patient characteristics, prior medical records, and ICU outcomes. Other hospitals $\{H_{i}\}$ each has their patient data: $\{\mathcal{D}_{i}\mid i\in[1..N]\}$.
For this binary prediction task, hospitals typically optimize for performance metrics, for example the area under the receiver-operating characteristic curve (AUC). Using only their data, $H_{o}$ can train a model $\mathcal{M}$ with parameters $\theta$ to achieve:
(Baseline Performance) $\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$
where $f$ is their chosen algorithm with parameter $\theta$.
When $H_{o}$ has exhausted their own internal data, they may benefit from incorporating additional target data sources $T\subset[1..N]$. By combining datasets, i.e., $\mathcal{D}_{T}=\{\mathcal{D}_{i}\mid i\in T\}\cup\mathcal{D}_{o}$, $H_{o}$ can potentially achieve better results:
(Combined Performance) $\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$
We define the potential improvement from data addition as $\delta_{T}=\delta_{(o,T)}=\mathrm{AUC}_{T}-\mathrm{AUC}_{o}$. To add a single additional data source by setting $T=\{i\}$, the improvement is $\delta_{i}=\delta_{(o,i)}=\mathrm{AUC}_{i}-\mathrm{AUC}_{o}$. This leads to our central question: Without seeing target data, how does a hospital ascertain potential data sources to combine with?
Formally, given $n\leq N$, we seek a strategy $\pi$ that selects $n$ target datasets $T=\pi(\mathcal{D}_{o},n)$ to maximize model utility:
(Ideal Dataset Combination) $\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset{[1...N]\choose n}}\mathrm{AUC}_{T}$
Practical Considerations.
Computing every subset $T\subset{[1...N]\choose n}$’s associated $\delta_{T}$ is exponential in $n$. To make this problem tractable, we make two key assumptions. First, we apply strategies greedily, selecting top-ranked target datasets. With the ultimate objective of improving the source hospital’s prediction task, we fix $H_{o}$; to compare the trade-offs between strategies in Section 3, we apply each $\pi$ greedily to select top-$n$ institution(s) for $H_{o}$ without replacement. Second, in in data constrained settings, we aim to maximize the probability of positive improvement: $P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)$.
Kullback–Leibler Divergence.
Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work (shen2024data, ). KL divergence (kullback1951information, ), also called information gain (quinlan1986induction, ), describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
(Kullback–Leibler Divergence) $\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$
Because computing KL-divergence on datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial,  (shen2024data, ) proposes two groups of scores to make this divergence approximation tractable from small samples. Specifically, score $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covariates — with the goal of inferring dataset membership. Then, the resulting model’s probability score function $\text{Score}(\cdot):\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
(KL-XY Score) $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$
Details are described in Section 3.
Privacy Model for $\pi_{p}$.
We operate under a semi-honest privacy model—also known as honest-but-curious or passive security—where parties follow protocols but may probe intermediate values. Parties are “curious”, meaning that they can probe into the intermediate values to avoid paying for the data. This assumes a weaker security model than malicious security where a corrupted party may input foul data, but ensures the algorithm to be private throughout the computation. This privacy preservation model incentivizes collaboration, improving upon methods in  (shen2024data, ) .
MPC Preliminary
To secure this divergence computation cryptographically, Secure Multiparty Computation (MPC) (yao1982protocols, ; shamir1979share, ) protocols are leveraged. Specifically, in $\mathrm{SecureKL}$, each party encodes $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ to preserve privacy for both parties. This is implemented with the research framework CrypTen (knott2021crypten, ), specialized for MPC and machine learning. Our algorithmic and engineering details are in Sections  3 and  4, respectively. For related secure techniques, see Section 5.4.
Additional Assumptions
Generally, we consider high stakes domains where disparate data may have additive benefits to the existing data. In order to make privacy boundaries tractable, we make the following additional assumptions:
  1. (1)
    Existing knowledge is not private. The hospitals are aware of each other having such data to begin with. The hospitals may know of the available underlying dataset size and format, which is assumed to be uniform across the hospitals in the setup to simulate unit-cost. Hospitals frequently know of each other’s resources, and the available ICU units are contentious, not kept secret.
  2. (2)
    Uniformity of $|\mathcal{D}_{i}|$. Though each hospital gets to price their data and set their own budget, for generality, the uniformity assumption allows us to use the number of additional data sources $n$ as the main ”budget proxy” across different strategies.
  3. (3)
    Legal risks of sharing any data are omnipresent in high stakes domains. The risks with sharing sensitive data in $\pi_{d}$ and $\pi_{s}$ are not made explicit, but assumed to be ”medium” and ”medium-to-high” respectively. This abstraction side-steps legal discussion, which would go beyond the scope of our paper.
  4. (4)
    No malice is assumed on any of the parties involved, as each hospital wants to authentically sell their data and set up a potential collaboration. This assumption becomes stronger when the number of parties grows or when the setup changes to potentially more competitive industries with less trust. We note our limitations in Section 4.5.
3. Methods
Figure 1. Our Method $\mathrm{SecureKL}(\mathcal{D}_{o},\mathcal{D}_{i})$. Each side encrypts their data, and a model is trained on their joint data. Using private KL-based measures, their distance is computed, and the final result is revealed after decryption, which requires both parties.
[…]
Data is encrypted on each side.
We summarize our data acquisition strategies differentiated by leakage risks, which correlate with potential costs:
Category 1, medium-to-high leakage, sharing raw data. $\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered (lds, ). Though leakage can be controlled through $k$, the data is inherently sensitive. The underlying distance uses  (shen2024data, ) ’s KL-XY Score.
Category 2, medium leakage, sharing summary statistics. $\pi_{d}(n)$ uses demographic metadata to guide data selection. This is implemented through ratio distance between source and target distributions, which may be considered aggregates therefore potentially not sensitive, such as when the underlying aggregation function $\phi$ is differentially private.
Category 3, zero leakage, sharing no additional information besides what is assumed public. There are two methods: a. Blind selection baseline: $\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. Prior works suggests that when $n=1$, randomly selecting a source in hospital ICU may be risky and inefficient. b. Our method $\pi_{p}(n)$ selects data sources based on privacy-preserving measure for data combination, specifically Private KL-XY.
3.1. Trivial Baseline: Blind Selection
Blind selection refers to the process when no information is provided. $\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. This random strategy may evade selection biases and help gather diverse data. Yet, prior work  (shen2024data, ) suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
3.2. Sharing Summary Statistics
A relaxation to sharing no sensitive data is to share metadata. While demographic traits are often causal and available, their exact cause in relation to the task is not a priori established (without a highly effective model), therefore their success in distributional-matching is not guaranteed to be strong. Additionally, in practice, the most effective model that results from data combination may or may not be causally-sound. Nevertheless, we posit alternative strategy $\pi_{d}(n)$ to find the demographically close candidates to guide data selection: Let $\phi:\mathcal{D}\to\mathbb{R}^{m}$ be an $m$-dimensional summary statistic of a demographic feature i.e. the racial distribution of patients. Then, we use the distributional distance between $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$, characterized by their $L_{2}$-distance through $\phi$:
(Demographic-based Strategy) $\pi_{d}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,L_{2}(\phi(\mathcal{D}_{o})||\phi(\mathcal{D}_{i})).$
3.3. KL-based Methods, in Plaintext
$\pi_{s}(n,k)$ assumes each of the candidate hospitals will share a set of raw data. In ICU data, simulate that a default of $1\%$ is shared, so $k=3000\times 1\%=30$, though we run experiments with $k\in\{3,30.300,3000\}$ (Section 4.3). Though leakage can be controlled through $k$, the data is inherently sensitive.
This is implemented with KL-based measures similar to (shen2024data, ) . To recap, $\mathrm{KL}(P||Q)$ is not symmetrical, meaning that it is not a ”metric” that satisfies triangle inequality. Intuitively, this means the measure is directional: a hospital’s distribution $P_{o}$ may be ”close” to the target distribution $P_{i}$, but not the other way around:
(Ideal Estimator) $\mathrm{KL}(P_{o}||P_{i})=\int_{x\in\mathcal{X}}\log\frac{P_{o}(\mathop{}\!\mathrm{d}x)}{P_{i}(\mathop{}\!\mathrm{d}x)}P_{i}(\mathop{}\!\mathrm{d}x)$
Because we only have access to finite data $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$, approximations are needed. Typically, a learned model can capture distributional information, used to estimate continuous entropy. Thus the joint distribution of features and labels from both the source and target are included, with the goal of deriving an efficient estimator for $\mathrm{KL}(P_{o}||P_{i})$ that captures distributional shift from source to target.
Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model (Cox1958TheRA, ) on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covaraiates – with the goal of inferring dataset membership. A score of $0.5$ or less means the datasets are not distinguishable, making the data potentially useful.  (shen2024data, ) established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task. We note again that even though this model is trained on both parties’ data, the final algorithm that the hospital uses to train on combined data is not restricted.
Then, the resulting model’s probability score function $\text{Score}(\cdot):\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
(KL-X) $\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}(x)).$
(KL-XY) $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$
We focus on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, and reproduce (shen2024data, ) ’s results that it is predictive of downstream change in AUC.
Let the score function $g_{\mathrm{KL}}$ be the approximate of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$. The strategy selects the most likely hospital with the closest distance under the measure:
(KL-based Strategy, in plaintext) $\pi_{s}(n=1,k=K)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{KL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$
When only a subset is available, this function is adjusted by swapping $\mathcal{D}_{i}$ for $\mathcal{D}_{i}^{\prime}\subseteq\mathcal{D}_{i}$ where $|\mathcal{D}_{i}^{\prime}|=k$. We denote the full dataset size as $K=|\mathcal{D}_{i}|$.
3.4. SecureKL: Private KL-based Method
Using MPC, we extend on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to require no information sharing (besides what was already assumed public). Specifically we leverage the MPC based framework provided by CrypTen (knott2021crypten, ), a library designed for privacy-preserving machine learning, to implement private $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. As illustrated in Figure 1, the logistic regression as well as the scoring need to be implemented in private. Our code is publicly available  [^1] https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
Denote the private encoding of $x$ as $[x]$.
(Secure KL-X) $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}([x])).$
(Secure KL-XY) $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}([x,y])).$
Let the score function $g_{\mathrm{SKL}}$ be the secure approximation of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$. The strategy selects the most likely hospital with the closest distance under the measure:
(SecureKL Strategy, encrypted) $\pi_{p}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{SKL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$
As shown in Figure 1, any KL-based measure $g_{\mathrm{SKL}}$ can be adapted to our setup. We mainly use $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ as the underlying measure. Its performance is detailed in Section 4.4. Additionally, even though our implementation measures distance of data between one source and one target party, the setup readily extends to accommodating multiple parties. We note the engineering limitations in Section 4.5.3.
4. Experiments
4.1. Experimental Setup
We validate our method and demonstrate its applicability using the eICU Collaborative Research Dataset (Pollard2018TheEC, ), which contains over $200,000$ admissions from $208$ hospitals across the United States. Following the data cleaning and exclusion criteria outlined by (Water2023YetAI, ) and (shen2024data, ), we selected the $12$ hospitals with the highest number of patient visits (each with at least $2000$ patients) as our $\mathbf{H}$. Each strategy would compute with the same $K=3000$ records, as the total available data per hospital.
We simulate the problem setup for each hospital with the $24$-hour mortality prediction task. The strategy comparisons described in Section 3 are implemented using $1500$ samples and the AUC is evaluated on $400$ samples for all of our experiments unless otherwise noted. This follows training and evaluation protocols in Yet Another ICU Benchmark (Water2023YetAI, ). For the data combination experiments that compute AUC change $\delta_{i}$ or $\delta_{T}$, to match  (shen2024data, ) , we take $1500$ random samples from each selected dataset and combine it with $1500$ samples from $\mathcal{D}_{o}$ (fixed across all experiments).
Implementing $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be privacy-preserving requires training logistic regression model in private. This is used in the private setting to estimate $g_{\mathrm{SKL}}$$\mathrm{Score}([X])$ or $\mathrm{Score}([X,Y])$ – for each pair of hospitals. Our experiments train encrypted logistic regression in CrypTen (knott2021crypten, ) using the library’s SGD optimizer. To ensure a fair comparison between the scores obtained through plaintext and encrypted settings, we re-implement plaintext Score($X$) and Score($X,Y$) using logistic regression with SGD in PyTorch (paszke2019pytorch, ). This is because encrypted version of L-BFGS – the optimizer prior work (shen2024data, ) uses in plaintext-only with Scikit-Learn (scikit-learn, )– is not available in CrypTen, though it leads to better downstream performance. Hyper-parameter tuning for SGD in private and plain text are performed independently, with the details in Appendix B.
4.2. Experimental Questions
We ask three sets of questions:
  1. (1)
    Consistency: Does using multiparty implementation sacrifice original measure’s effectiveness? Practically, we evaluate this through the analysis of private and plaintext scores. For our selected metrics, we expect AUC change to be negatively with KL-based measures—meaning the closer the additional target dataset is to the source hospital, the more the AUC will improve compared to other potential target datasets. Section 4.3 tests the correlation of our private scores and plaintext scores with full access (setting $k=K$). In addition to computing Spearman’s rank correlation coefficient of the KL-scores, we probe the discrepancy of the downstream effect between $\{\delta_{i}|i=\pi_{s}(n=1,k=K)\}$ using plaintext $\mathrm{KL}_{\mathcal{X}},\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ with all the underlying data and $\{\delta_{i}|\pi=\pi_{p}\}$ using $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ for each source hospital $H_{o}\in\mathbf{H}$.
  2. (2)
    Positivity: Does our method pick hospitals that reliably improve performance? If source dataset $D_{o}$ can only add data from $n$ more hospitals, does our measure lead to eventual AUC improvements? In Section 4.4, we test our framework on a multi-dataset combination experiment and find that it successful improves the source hospital’s downstream outcome. Specifically, when selecting a single additional data source ($n=1$), all but 2 hospitals improves, and when selecting top 2 or 3, all hospitals see a positive AUC change. This shows a consistent added benefit.
    Lastly, Section 4.4.1 compare with different strategies proposed in Section 3, and Section 4.4.2 analyze the benefits of using private dataset combination $\mathrm{SKL}$.
  3. (3)
    Error analysis: If our privacy-preserving method is not the dominant strategy against alternatives including limited data accessibility, why is that? Section 4.5 performs additional analyses on (a) hospitals with low $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ correlations, and (b) hospitals lagging AUC improvements using the random strategy $\pi_{0}$ or limited-sample strategy $\pi_{s}$ for selecting $n=3$ candidate hospitals.
Lastly, we note engineering hurdles to scale to the real-world in Section 4.5.3. We hereby detail our results.
4.3. Consistency Between Plaintext and Encrypted Computations
Our encrypted computations are programmed with CrypTen, when the plaintext counterpart is PyTorch-only. We show using $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ lead to highly comparable behavior as $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}}$, respectively.
Spearman’s Rank Correlation Coefficient for Underlying Scores
For each source hospital $H_{o}$, use all full samples for $\mathcal{D}_{i}$. Between $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ on $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ for all remaining hospitals $H_{i}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.908$ with a range of $[0.691,1.0]$, obtaining $p<0.02$ across all hospitals. Between $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.9303$ with a range of $[0.455,0.991]$, with 11 of 12 hospitals achieving p-values below $0.05$. After applying Hochberg false discovery rate correction (Benjamini1995ControllingTF, ), our p-values remain significant. This range is an artifact of sweeping hyperparameters independently in plaintext and encrypted optimisations. When we unify the SGD hyperparameters, we indeed get a tighter range. For all 12 hospitals, see appended Appendix C for details.
Downstream Model Improvements
We further simulate the effect by adding encryption through its impact on the downstream AUC. This examines whether there will be a shift in the full hospital ranking, if we switch from a plaintext setup to encrypted. For $H_{o}\sim\mathbf{H}$, we measure $\delta_{i}$ that results from adding $\mathcal{D}_{i}$ to $\mathcal{D}_{o}$ for all $i$. This correlates all target hospitals $\{H_{i}\}$ with their ground truths $\{\delta_{i}\}$ in the case of picking a single target hospital. We find the linear coefficient for encrypted $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.182$ and plaintext $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.184$ ($99\%$ matching). Both $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$ have a linear coefficient of $-0.164$ with $\delta_{i}$. For all strategies’ correlations with ground truth at $n=1$, see Appendix A.
Figure 2. AUC change $\delta_{T}$, after including top-$3$ hospital per strategy, over all source hospitals. Our private strategy $\pi_{p}$ is compared with (a) demographic-based $\pi_{d}$ for gender, age, and race. (b) plain text limited-sample $\pi_{s}(k=K)$ for $k=300$ (10%) and $k=30$ (1%).
4.4. Positivity of SecureKL and Practical Implications
Overall Positivity
We evaluate the practical utility of SecureKL by applying it in a multi-source data combination experiment, where $n\in\{1,2,3\}$. For $n=1$, we find that $\pi_{p}$ improves AUC in 10 out of the 12 hospitals. When $n=2$ and $n=3$, we find that using $\pi_{p}$ consistently improves AUC for all hospitals. Overall, 34 out of the 36 dataset combinations we evaluate on have an AUC improvement $\delta_{T}>0$, suggesting that $\pi_{p}$ is a reasonable strategy for selecting hospital dataset combinations with a high expected return $\mathbb{E}[P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)]$ for the source hospital from using our strategy.
4.4.1. Comparing With Alternative Strategies
Other strategies – $\pi_{0}$, $\pi_{d}$, and $\pi_{s}$ – can also arrive at positive datasets. Comparing private method $\pi_{p}(n=3)$ to other strategies at $n=3$, we find the following results, illustrated in Figure 2:
  1. (1)
[…]
  2. (2)
    Demographic-based strategies underperform compared to $\pi_{p}$ on average. However, we observe that $\pi_{d}$-gender can be highly effective for a subset of hospitals, as it achieves the highest 75th percentile (Q3) of $0.033$ among all strategies. This indicates that for 25% of hospitals, $\delta_{T}\geq 0.033$. Despite this, $\pi_{d}$-gender has a lower median value of $0.012$ compared to $\pi_{p}$, exhibits a high standard deviation ($0.022$), and degrades the performance for certain hospitals. Similarly, $\pi_{d}$-age has a median of $0.014$, and $\pi_{d}$-race has a median of $0.008$, both lower than $\pi_{p}$’s median.
  3. (3)
    Plaintext small-sample strategies, $\pi_{s}$, outperform all demographic-based methods but slightly underperform relative to $\pi_{p}$. For instance, $\pi_{s}(k=300)$ has a median $\delta_{T}$ of $0.0178$, and although it achieves $\delta_{T}>0$ across all hospitals, it performs worse on average compared to $\pi_{p}$ and exhibits a higher standard deviation ($0.017$). $\pi_{s}(k=30)$ has a median $\delta_{T}$ of $0.0165$. Compared to other strategies, it has the largest standard deviation ($0.024$), and it degrades the performance for some hospitals.
In summary, our method $\pi_{p}$ achieves the highest AUC improvement on average with the lowest standard deviation, demonstrating consistent improvement for all hospitals. In contrast, demographic-based and plaintext small-sample strategies exhibit greater variability, with some strategies improving performance for specific subsets of hospitals but underperforming or degrading results in others.
4.4.2. SecureKL Analysis
After establishing that $\pi_{p}$ with $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$is a robust strategy in practical downstream performance, we hereby synopsize the benefits of $\mathrm{SecureKL}$ and elaborate on their practical implications.
A Principled Approach To Data Minimization. Our major contribution is to match plaintext performance with no data sharing. Using MPC provides input privacy, meaning that if both hospitals only want to know the resulting score, the computation can be done without leaking original data. This strong guarantee can significantly ease the tension related to privacy and compliance in setting up a collaboration, leading to a practical ”data appraisal stage” in data-limited high stakes domains.
In the case where that output can be sensitive, i.e., when hospitals query each other multiple times and accrue information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response (dwork2014algorithmic, ).
In theory, any data combination method (if Turing-complete) can be made private; yet, in practice, balancing the right trade-off of utility and privacy is non-trivial. Barring engineering difficulties, not all algorithms readily adapt efficiently in private. Prior work  (xu2022data, ) included the trained model and test data in private; while relatively exact, complex methods would exacerbate the same operational limitations discussed in Section 4.5.3.
Gain from Data Availability. In contrast to limited-sample approaches, a key advantage for our method $\pi_{p}$ is that it takes advantage of all of the underlying data – generally impossible with non-secure methods for private data in heavily regulated domains. The general intuition is that data is localized; therefore, once a good target hospital is identified, we should acquire all of the data. It may be tempting to assert that we prefer the highest $k$ for data addition algorithms as well. In our experiments, while this is generally true, the smaller $k$ sometimes outperform larger $k$ in plaintext strategy $\pi_{s}$, which we investigate in Section 4.5 and in Figure 5. This occasionally non-monotonic behavior mirrors the challenge of data combination itself: even within one source dataset for the same estimator, more data is not necessarily better. This suggests a domain-specific alternative to sharing a large amount of data for some source hospital, and points to future directions to using secure computation on a minimal-sized sample dataset for minimal performance overhead while remaining private.
4.5. Error Analysis
4.5.1. Underlying Score Limitations
Data addition algorithms underpin the effectiveness of our method. Even if $\mathcal{D}_{o}$ obtains access to all the plaintext data, there is no guarantee that $\pi_{p}$ can correctly predict whether the data is useful. As seen in Figure 4, Hospital 243’s utility when acquiring another data set is badly correlated with plaintext and encrypted KL-XY scores. This leads to its bad strategy for acquiring the top 3 hospitals, as seen in the middle pane of Figure 5. Interestingly, for this hospital, no other informational strategy excels, either, so choosing a random 3 may be preferred.
This behavior stems from the underlying measure, not from adding secure computation: in Figure 4 and Figure 5, the encrypted performance closely follows that of plaintext performance, for both good and bad downstream correlations.
4.5.2. Sometimes, Not All Underlying Data Is Needed
Relatedly, when seeing a few samples can successfully identify useful candidate hospitals, $\pi_{p}$ does not always outperform $\pi_{s}$ on small samples.
In the right panel of Figure 5, hospital 199, the smaller sample sizes achieve a score that better reflects ground truth as a data addition strategy. In that case, the hospital may not need the full sample to know which target hospitals to collaborate with.
This behavior is specific to the interaction of the data and the underlying score, and does not affect the general insight that adding private computation preserves privacy (and eases privacy-related risks that hinder data sharing). We further note that our method still clearly applies to encrypted computation on a smaller data set under data minimization.
Figure 3. Hospital 243: Underlying KL-score performs poorly
[…]
NEED UPDATING some conclusion and description
Figure 4. Hospital 420: Underlying KL-score performs well
[…]
NEED UPDATING
Figure 5. Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{p}(k=30)$ and $\pi_{0}$. Middle: All strategies perform similarly. Right: $\pi_{p}(k=300)$ outperforms $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. In all panels, the bars represent the standard deviation.
[…]
NEED UPDATING
4.5.3. Encrypted Computation Limitations
Engineering a secure system for machine learning requires both machine learning and software engineering knowledge. We share our our code and method, but also note potential challenges with deploying our secure computation:
  1. (1)
    Operational: engineering personnel limitations. While our implementation requires little cryptographic knowledge to deploy, it still needs technically-trained staff at each participating hospital to collaborate and maintain.
  2. (2)
    Engineering: Extending any MPC protocol is non-trivial, as security engineering is a specialized skill. While $\mathrm{SecureKL}$ applies broadly to other underlying scores in multi-party setups, every new algorithm requires software engineering - prototyping, tuning, debugging —which can be especially costly for hospitals.
  3. (3)
    Framework Limitation: While CrypTen is designed to accommodate PyTorch, it is a research tool where not all plain text functionalities are implemented. Writing optimizers – such as L-BFGS – and custom operators that are not readily available requires both machine learning and cryptography knowledge.
  4. (4)
    Inherent to Secure Computation: When the method requires significant hyper-parameter tuning, such as using SGD on small batch data with learning rate schedules, plaintext tuning may not transfer perfectly. As detailed in Appendix  B, our hyperparameters for SGD are indeed different in encrypted and plaintext settings. However, encrypted computation hides loss curves and training details by default, complicating development.
5. Related Work
5.1. Data Valuation and Pricing
The question of how to assess the impact and therefore worth of data has been well studied. Data valuation as a field seeks to quantify the contribution of individual data points or datasets to model performance. Shapley value-based approaches provide theoretically grounded valuations but scale poorly to large datasets (ghorbani2019data, ). More efficient methods include influence functions (koh2017understanding, ) and leave-one-out testing (cook1979influential, ). Recent work has extended these concepts to dataset-level valuation (jia2019towards, ). As practitioners create larger datasets, the emergence of data marketplaces has sparked interest in data pricing mechanisms (kumar2020marketplace, ). Query-based pricing (koutris2015query, ) and outcome-driven valuations (radic2024pricing, ) aim to balance seller compensation with buyer utility. While these approaches inform fair data exchange, they typically assume direct access to data, unlike our privacy-preserving method.
5.2. Alternative Approaches to Data Sharing
Recent work has explored several approaches to mitigate data sharing constraints while maintaining model performance. We discuss two primary directions: synthetic data generation and transfer learning from public pretraining.
Synthetic data generation
Synthetic data generation has emerged as a promising approach to expand training datasets while preserving privacy. Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data (goncalves2020generation, ), medical imaging data (thambawita2022singan, ), and electronic health records (baowaly2019synthesizing, ). These methods can preserve statistical properties of the original data while providing differential privacy guarantees. Transforming data into a similar form that de-sensitizes certain attributes can be desirable (drechsler2011synthetic, ; howe2017synthetic, ; nikolenko2021synthetic, ; gonzales2023synthetic, ; sweeney2002k, ). Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself (jordon2021hide, ). Additionally, outside of the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks (narayanan2006break, ; jordon2021hide, ).
However, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables (goncalves2020generation, ). For tabular data, methods like CTGAN (xu2019modeling, ) and TVAE (xu2019modeling, ) have demonstrated ability to learn complex distributions while preserving correlations between features. However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance (assefa2020generating, ). Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees (jordon2018pate, ). While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset (yang2024tabular, ).
Public pre-training and private fine-tuning
Transfer learning via public pretraining has become increasingly popular for domains with limited private data access. BioBERT (lee2020biobert, ) and ClinicalBERT (alsentzer2019publicly, ) demonstrated that pretraining on PubMed abstracts and clinical notes can improve performance on downstream medical tasks. Similar approaches have emerged in other regulated domains, including FinBERT (liu2021finbert, ) for financial applications. However, the effectiveness of transfer learning depends heavily on domain alignment. One study showed that continued pretraining on domain-specific data significantly outperforms generic pretraining when domains differ substantially (gururangan2020don, ). This presents challenges for highly specialized fields where public data may not capture domain-specific patterns (le2022language, ). Recent work has explored methods to quantify and optimize domain adaptation. Adaptive pretraining strategies (howard2018universal, ) and domain-specific vocabulary augmentation (gururangan2020don, ) have shown promise in bridging domain gaps. However, these approaches still require substantial compute and may not fully capture specialized domain knowledge present in private datasets.
5.3. Benefits of Data Scaling Beyond Performance
Recent work has demonstrated that increasing dataset size and diversity yields benefits beyond raw performance metrics. Large-scale training data has been shown to improve model robustness to distribution shifts (taori2020measuring, ) and reduce demographic performance disparities (chen2018my, ). Studies of vision models trained on increasingly large datasets show improved out-of-distribution generalization (miller2021accuracy, ). Similarly, language models trained on diverse data demonstrate better performance across different domains and demographic groups (dodge2021documenting, ).
5.4. Secure and Confidential Computation
Secure and confidential computation encompasses cryptographic techniques that protect information privacy during computation. In a two-party setup between a model owner and data owner, these methods enable computing joint functions on private inputs without revealing them to other parties.
This requires an encoding scheme $\mathrm{Enc}(\cdot)$ that satisfies the homomorphic property: $\mathrm{Enc}(A)\circ\mathrm{Enc}(B)=\mathrm{Enc}(A\circ B)$, where $A$ and $B$ represent data held by two parties. The inverse function $\mathrm{Enc}^{-1}(\cdot)$ must exist to decode the final output: $\mathrm{Enc}^{-1}(A\circ B)=A\circ B$. Considering an “honest-but-curious” threat model, where parties aim to jointly compute on privately-held data, two main approaches emerge.
Fully Homomorphic Encryption (FHE)
FHE enables arbitrary additions and multiplications on encrypted inputs. While it represents the gold standard for encrypted computation, adapting it to modern machine learning is challenging due to computational constraints from growing ciphertext size. FHE implementations typically use lattice-based schemes requiring periodic “bootstrapping” (key refreshing and noise reduction through re-encryption) via methods like the CKKS scheme (cheon2017homomorphic, ). This introduces cryptographic parameters that non-experts struggle to configure effectively.
Secure Multi-party Computation (S-MPC)
SMPC enables multiple parties to compute functions over private inputs while revealing only the final output (yao1982protocols, ; shamir1979share, ). An MPC system uses key exchanges, encryption schemes, and secure communication to ensure only encrypted data leaves owner control (micali1987play, ). Though generally faster than FHE, SMPC’s engineering complexity and communication overhead can limit adoption. Traditional private training approaches that completely hide data can also impede essential model development tasks like inspection, monitoring, and debugging.
5.5. Secure Data Combination
Recent work has explored methods for securely combining datasets while preserving privacy and improving model performance. Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data (aono2017privacy, ). However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets (mcmahan2017communication, ). More recent techniques have introduced frameworks for evaluating potential data partnerships before commitment. These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets (leung2019towards, ; chakraborty2024privacy, ). Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information (duan2021flexible, ). Several systems have been developed to facilitate secure data combination in specific domains. In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance (raisaro2018m, ). Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality (liu2021finbert, ).
5.6. Other Privacy-Preserving Methods
Federated Learning.
Cross-silo federated, decentralized, and collaborative machine learning  (mcmahan2017communication, ; li2020federated, ; bonawitztowards, ; kairouz2021advances, ) focus on acquiring more data through improved data governance and efficient system design. Healthcare machine learning is considered especially suitable, as health records are often isolated (rieke2020future, ; xu2021federated, ; nguyen2022federated, ). Yet, even though no raw data is shared, model parameters or gradients flow through the system. As the federated computing paradigm offer no privacy guarantee, the system is vulnerable to model inversion (geiping2020inverting, ) and gradients leakage attacks (boenisch2023curious, ; zhu2019deep, ). A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimisation is traded off with privacy (lyu2022privacy, ; raynal2024conflict, ). Building on the insight that useful data is often disparately owned, we tackle the specific incentive problem between pairs of data players where one side trains the model, instead of scaling up a federation (number of parties) to address data access issues. We thus focusing on making this exchange efficient, accurate, and private.
Compared to vanilla Federated Learning, an MPC system (shamir1979share, ; yao1982protocols, ; bonawitz2017practical, ; knott2021crypten, ) provides stronger guarantee in terms of input security. Model owners and data owners can potentially federate their proprietary data, including model weights, training, and testing data, can work together under stringent privacy requirements. Our work extends the line of works  (xu2022data, ; yang2024fedfed, ; bonawitz2017practical, ) that demonstrates the potential of incorporating MPC in various federated scenarios. On the practical side, unlike mobile-based networks for secure federated learning protocols (bonawitztowards, ), our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
Differential Privacy
Differential privacy (DP) (dwork2006differential, ) offers formal privacy guarantees for sharing data and training machine learning models. While DP mechanisms can protect individual privacy when releasing model outputs or aggregated statistics, they face significant limitations for inter-organizational data sharing. The primary challenge is that DP operates on already-pooled data, but organizations are often unwilling to share their raw data in the first place (dwork2014algorithmic, ). Even when organizations are willing to share data, the privacy guarantees of DP come at a substantial cost to utility, particularly in machine learning applications. DP-SGD, the standard approach for training deep neural networks with differential privacy, significantly degrades model performance compared to non-private training (abadi2016deep, ). This performance impact is especially pronounced in data-constrained settings, where recent work has shown that large models rely heavily on memorization of rare examples that DP mechanisms tend to obscure (feldman2020neural, ). The privacy-utility trade-off becomes even more challenging when dealing with high-dimensional data or complex learning tasks. Studies have demonstrated that achieving meaningful privacy guarantees while maintaining acceptable model performance requires prohibitively large datasets (bagdasaryan2020backdoor, ). This limitation is particularly problematic in specialized domains like healthcare, where data is inherently limited and performance requirements are stringent (geyer2017differentially, ). Recent work has attempted to improve the privacy-utility trade-off through advanced composition theorems and adaptive privacy budget allocation (papernot2021tempered, ). However, these approaches still struggle to match the performance of non-private training, especially when working with modern deep learning architectures (tramer2020differentially, ). While differential privacy offers important theoretical guarantees, our work focuses on the practical challenge of enabling data owners to evaluate potential partnerships before sharing any data, addressing a key barrier to collaboration that DP alone cannot solve.
6. Conclusion
Our work demonstrates that privacy-preserving data valuation can help organizations identify beneficial data partnerships while maintaining data sovereignty. Through SecureKL, we show that entities can make informed decisions about data sharing without compromising privacy or requiring complete dataset access. As the AI community continues to grapple with data access challenges, particularly in regulated domains like healthcare, methods that balance privacy and utility will become increasingly critical for responsible advancement of the field. As noted in Section 4.5, our approach has several limitations, including the fact that, despite impressive aggregate results, our method is less effective for individual hospitals; this finding is fertile ground for future work. Additionally, our work present opportunities for follow-up research. Our method assumes static datasets and may not generalize well to scenarios where data distributions evolve rapidly over time. A sequential version of our framework may more closely model dynamic data collaborations. Future work should explore extending these techniques to handle more complex data types and dynamic distribution shifts while maintaining strong privacy guarantees.
Appendix A Correlation with downstream performance
On Table 1, we report the Pearson correlations between $\pi_{s}(k=K)$ for $k\in\{3,30,300,3000\}$ and $\delta_{i}$. On Table 2, we report the Pearson correlations between different strategies and $\delta_{i}$.
k $\rho$ p-value
3 -0.063 4.70e-01
30 -0.082 3.47e-01
300 -0.059 5.00e-01
3000 -0.184 3.47e-02
k $\rho$ p-value
3 -0.158 7.02e-02
30 0.167 5.60e-02
300 -0.097 2.70e-01
3000 -0.284 9.47e-04
Table 1. $\rho$ and p-value between AUC drop and plaintext KL using k samples using SGD (left) and LBFGS (right).
Data Addition $\pi$ $r$ p-value
$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$0.182 3.65e-02
$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$0.162 6.27e-02
$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$0.184 3.47e-02
$\mathrm{KL}_{\mathcal{X}}$0.162 7.13e-02
Gender 0.097 2.65e-01
Race 0.018 8.29e-01
Age 0.053 5.33e-01
Table 2. Using the eICU dataset, we measure the Pearson correlation $r$ between the strategy $\pi$ and data addition AUC drop. p-values below .05.05.05.05 are bolded.
Appendix B Hyperparamter Tuning
We obtain Score(X,Y) by training a Logistic Regression model using SGD. We find that SGD requires hyper-parameter tuning in order to perform well when evaluated on Brier Score Loss. We used Optuna to perform hyper-parameters search. The hyper-parameters we use for plaintext scores are:
  1. (1)
    learning rate: $0.0795$
  2. (2)
    patience: 2
  3. (3)
    tolerance: 0.000117
  4. (4)
    momentum: 0.886
  5. (5)
    weight decay: 1.81e-9
  6. (6)
    dampening: .0545
The hyper-parameters for the encrypted model:
  1. (1)
    learning rate: $0.0974$
  2. (2)
    patience: 5
  3. (3)
    tolerance: 0.000132
  4. (4)
    momentum: 0.907
  5. (5)
    weight decay: 8.14e-7
  6. (6)
    dampening: .0545
Appendix C Correlations between Encrypted Scores and Plaintext Scores
Hospital $\rho(\mathrm{KL}_{\mathcal{X}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}})$ p-value $\rho(\mathrm{KL}_{\mathcal{X}\mathcal{Y}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}})$ p-value
73 0.945 1.118e-05 1.000 0.0
264 0.973 5.142e-07 0.945 1.118e-05
420 0.982 8.403e-08 0.991 3.763e-09
243 0.973 5.142e-07 0.909 1.056e-04
338 0.973 5.142e-07 0.982 8.403e-08
443 0.964 1.852e-06 0.882 3.302e-04
199 0.991 3.763e-09 0.973 5.142e-07
458 0.873 4.546e-04 0.964 1.852e-06
300 0.455 1.601e-01 0.691 1.857e-02
188 0.718 1.280e-02 0.864 6.117e-04
252 0.873 4.546e-04 0.809 2.559e-03
167 0.764 6.233e-03 0.891 2.335e-04
Table 3. Spearman Correlations $\rho$ for encrypted (in CrypTen) and plaintext (in PyTorch) KL-based methods
On Table 3, we measure the Spearman correlations between $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$, and between $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ for all hospitals. We find that all hospitals have statistically significant correlations with the exception of hospital 300’s $\rho$($\mathrm{KL}_{\mathcal{X}}$,$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$)