### v3

# Privacy-Preserving Dataset Combination

Abstract
Privacy concerns and competitive interests impede data access for machine learning, due to the inability to privately assess external data’s utility. This dynamic disadvantages smaller organizations that lack resources to aggressively pursue data-sharing agreements. In data-limited scenarios, not all external data is beneficial, and collaborations suffer especially in heavily regulated domains: metrics that aim to assess external data given a source e.g., approximating their KL-divergence, require accessing samples from both entities pre-collaboration, hence violating privacy (Compton et al., 2023; Shen et al., 2024). This conundrum disempowers legitimate data-sharing, leading to a false “privacy-utility trade-off”. To resolve privacy and uncertainty tensions simultaneously, we introduce SecureKL, the first secure protocol for dataset-to-dataset evaluations with zero privacy leakage, designed to be applied preceding data sharing. SecureKL evaluates a source dataset against candidates, performing dataset divergence metrics internally with private computations, all without assuming downstream models. On real-world data, SecureKL achieves high consistency ($90\%$ correlation with non-private counterparts) and successfully identifies beneficial data collaborations in highly-heterogeneous domains (ICU mortality prediction across hospitals and income prediction across states). Our results highlight that secure computation maximizes data utilization, outperforming privacy-agnostic utility assessments that leak information.
1 Introduction
Figure 1: Privacy can dis-incentivize data collaborations. Without seeing external data, an organization has two strategies: i. blind default $\pi_{0}$: randomly selecting partnerships causes hesitation and hinders partnerships. ii. private evaluation $\pi_{p}$: securely assessing datasets before commitment.
Since Hestness et al. (2017), empirical works have predetermined data as a key driver to performance gains, via the so-called “scaling laws” (Kaplan et al., 2020; Brown et al., 2020). Yet, accessing and combining datasets is persistently challenging. As datasets have evolved from small to larger (Sun et al., 2017), more diverse (Zhang et al., 2020), and more compute-optimal (Hoffmann et al., 2022), the field of machine learning continues to seek more data  (Mahajan et al., 2018; Brown et al., 2020; Park et al., 2019; Reed et al., 2022; Sheller et al., 2020) and better ways to combine it  (Shen et al., 2023; Nguyen et al., 2023; Petty et al., 2024; Li et al., 2024; McKinney et al., 2020).
Strategically combining data from different sources promises enhanced models, but disempowers smaller organizations. While diverse, high quality data often improves performance, robustness, and fairness (Miller et al., 2021), access to such data significantly varies across entities and domains (Tenopir et al., 2011; L’Heureux et al., 2017). As domain-specific data becomes increasingly valuable (Alsentzer et al., 2019; Gururangan et al., 2020; Lee et al., 2020), data-owning entities are more reluctant to share it for free, opting instead to sell it in emerging data markets (Acemoglu et al., 2022; Huang et al., 2021; Liang et al., 2018). This dynamic disproportionately handicaps smaller organizations who lack both the resources to purchase data and the leverage to negotiate favorable sharing agreements.
Organizations may hesitate to commit to a potential partnership when unsure about the benefits. As Figure 1 illustrates, this “commitment issue” is not solely a privacy issue; it’s the inability to privately assess an external dataset’s utility before partnerships. This evaluation is, however, nontrivial. For example, adding in-domain data does not necessarily result in more performant models. Due to the precarious nature of domain shifts, machine learning models are sensitive to additional training data, often unpredictably when the source dataset is small (e.g., a single hospital’s data) – a phenomenon known as “the dataset combination problem” (Koh et al., 2021; Wang et al., 2022; Taori et al., 2020; Bradley et al., 2021; Meng, 2018; Shen et al., 2024).
Figure 2: The Dataset Combination Problem. Real-world data collaborations are inherently uncertain. AUC change for a source entity, after incorporating external data across hospitals and states. Each entry representas training on a pairwise data combination for a test entity (x-axis) Each column represents training on a data combination for a fixed test dataset, formed by pairing the fixed entity with an external dataset in the y-axis. Left: In eICU (Pollard et al., 2018), 10 out of 12 hospitals may see their mortality prediction model degrade for some potential hospital partners. Right: In Folktables (Ding et al., 2021), combining with random state leads to worse income prediction in 10 out of 15 states. (red is bad; exact values are reported in Appendix 15)
Ideally, all target data should be considered to reduce uncertainty in costly data collaborations. Yet, datasets owned by separate entities cannot be directly and fully accessed, significantly limiting the practicality of non-private dataset measures (Shen et al., 2024; Ilyas et al., 2022).
Our work directly addresses this crux by recognizing both privacy and competitive incentives. First, before committing to acquiring unseen data, we enable organizations to privately gauge the relative utility of candidate datasets. Second, we provide strong privacy guarantees required of entities operating under stringent regulations e.g. healthcare providers to navigate data acquisition.
Developers are often uncertain about the most effective model before more data becomes available. This renders a secure data appraisal stage by Xu et al. (2022), which requires model parameters, not applicable. In this opportunistic setting, we ask:
Can we ascertain the differential utility of prospective datasets, without knowing the final model?
We introduce SecureKL, the first divergence-based dataset measure with zero privacy leakage. Our key insight is that private divergence computations (via secure multiparty computation (MPC)) are more data-efficient than sharing samples, thus achieving high performance. By preserving performance, SecureKL presents a compelling guarantee: for both parties, privacy is fully protected while data utilization is maximized.
Contributions
A novel secure dataset-to-dataset evaluation protocol SecureKL (SKL) that reduces uncertainty in data utility under limited data and budget, producing privacy-preserving measures while using the maximum available samples. SKL achieves a $90\%$ correlation with privacy-violating counterparts across two real-world heterogeneous domains. Empirically, on ICU mortality prediction, SKL reliably chooses beneficial hospital(s) to partner with, outperforming data-leaking alternatives, including using demographic summaries or sharing data subsets.
Impact
We provide a practical solution for organizations seeking to leverage collective data resources while maintaining privacy and competitive advantages. Our major advantage lies in reliability, especially when small organizations cannot afford to invest in detrimental partnerships. These results demonstrate the potential for wider data collaboration to advance machine learning applications in high-stakes domains while promoting more equitable access to data. Our code is available, and can be readily deployed to demonstrate potential data value preceding collaborations [^1] https://github.com/kere-nel/secure-data-eval.
2 Background
2.1 Dataset Combination Problem
In high-stakes domains, incorporating additional datasets may degrade the model. In healthcare scenarios, both Compton et al. (2023) and Shen et al. (2024) showed that blindly acquiring new datasets can degrade model performance. We extend these findings by replicating the effect under our own experimental setup. As shown in Figure 2, combining data from one hospital or state with an external source often leads to inconsistent, and sometimes negative, changes in model performance. This non-monotonic behavior highlights the need to evaluate data partners before embarking on a full-fledged collaboration.
2.2 Trading Off Sample Utility For “Privacy”
Figure 3: Non-private evaluation strategies. iii. sub sampling $\pi_{s}$: a subset of the target’s data is shared. iv. demographic summaries $\pi_{d}$: the target entity discloses distributions by protected attributes, i.e. age, gender, or race.
Before an agreement, informed negotiations become impossible when entities do not expose data. In practice, data owners frequently resort to tiny samples (Centers for Medicare & Medicaid Services, 2025) or summary statistics (e.g., race, gender, age) for making decisions (Figure 3). Yet, in data-limited settings, model performance is highly sensitive to new input. These heuristics are fickle, as sparse traits or limited samples often fail to capture the entire dataset’s nuance and complexity, especially in heterogeneous domains, creating a perceived privacy-utility trade-off. We will show that secure computation, on the other hand, can avoid this tradeoff by measuring over entire datasets while maintaining privacy.
3 Problem Setup
Consider a binary prediction task for ICU patient mortality based on electronic medical records. A source hospital $H_{o}$ has historical patient data $\mathcal{D}_{o}$ containing static past patient characteristics, prior medical records, and ICU outcomes. Other hospitals $\{H_{i}\}$ each has their patient data: $\{\mathcal{D}_{i}\mid i\in[1..N]\}$.
For this binary prediction task, hospitals typically optimize for performance metrics, for example the area under the receiver-operating characteristic curve (AUC). Using only their data, $H_{o}$ can train a model $\mathcal{M}$ with parameters $\theta$ to achieve:
$\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$ (Baseline AUC)
where $f$ is their chosen algorithm with parameter $\theta$.
When $H_{o}$ has exhausted their own internal data, they may benefit from incorporating additional target data sources $T\subset[1..N]$. By combining datasets, i.e., $\mathcal{D}_{T}=\{\mathcal{D}_{i}\mid i\in T\}\cup\mathcal{D}_{o}$, $H_{o}$ can potentially achieve better results:
$\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$ (Combined AUC)
We define the potential improvement from data addition as $\delta_{T}=\delta_{(o,T)}=\mathrm{AUC}_{T}-\mathrm{AUC}_{o}$. To add a single additional data source by setting $T=\{i\}$, the improvement is $\delta_{i}=\delta_{(o,i)}=\mathrm{AUC}_{i}-\mathrm{AUC}_{o}$. This leads to our central question:
Without seeing target data, how does a hospital ascertain potential data sources to combine with?
Formally, given $n\leq N$, we seek a strategy $\pi$ that selects $n$ target datasets $T=\pi(\mathcal{D}_{o},n)$ to maximize model utility:
$\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset\binom{[1...N]}{n}}\mathrm{AUC}_{T}$ (Ideal Combination)
Practical Considerations.
Computing every subset $T\subset\binom{[1...N]}{n}$’s associated $\delta_{T}$ is exponential in $n$. To make this problem tractable, we make two key assumptions. First, we apply strategies greedily, selecting top-ranked target datasets. With the ultimate objective of improving the source hospital’s prediction task, we fix $H_{o}$; to compare the trade-offs between strategies in Section 4, we apply each $\pi$ greedily to select top-$n$ institution(s) for $H_{o}$ without replacement. Second, in data constrained settings, we aim to maximize the probability of positive improvement: $P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)$.
Kullback–Leibler Divergence.
Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work (Shen et al., 2024). KL divergence (Kullback and Leibler, 1951), also called information gain (Quinlan, 1986), describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
$\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$ (KL-Divergence)
Because computing KL-divergence on hospital datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial due to the high dimensionality of the data, Shen et al. (2024) proposes two groups of scores to make this divergence approximation tractable from small samples. We adopt one such score, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, which trains a classifier to distinguish between source and target samples and uses the classifier’s output probabilities to estimate dataset divergence. We describe the $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score in detail in Section 4.2.
Privacy Model
We operate under a semi-honest privacy model—also known as honest-but-curious or passive security—where parties follow protocols but may probe intermediate values. Parties are “curious”, meaning that they can probe into the intermediate values to avoid paying for the data. This assumes a weaker security model than malicious security where a corrupted party may input foul data, but ensures the algorithm to be private throughout the computation. This privacy preservation model incentivizes collaboration, improving upon Shen et al. (2024) and  Ilyas et al. (2022).
Dataset Divergence and Downstream Utility
Building on the finding of Miller et al. (2021) where a model’s in-distribution performance is related to its out-of-distribution performance (across all model choices), we confer that dataset divergence does predict downstream model’s performance after combining the data (Table 3), using metrics introduced in Shen et al. (2024) (Section 4.2). Intuitively, in-distribution quality is paramount in low-data settings, where dataset divergence can capture greater complexity and nuance than accessing a few traits.Yet, privacy is unresolved: divergence measures entail accessing both entities’ data (Shen et al., 2024; Ilyas et al., 2022), posing significant risks for heavily-regulated entities who are liable for any data exposure (Centers for Medicare & Medicaid Services, 1996; Gervasi et al., 2022).
Secure Multiparty Computation (MPC)
To cryptographically secure our divergence computation, we use Secure Multiparty Computation (MPC)  (Yao, 1982; Shamir, 1979). MPC lets two or more parties to compute a function over private inputs, revealing only the final output (Goldwasser and Micali, 2019).
Engineering machine learning workflows in private faces differs drastically from non-private machine learning engineering, due to 1. precision configurations and normalizations (to avoid blow-ups and underflows), 2. error control and performance engineering for machine learning in bit-limited spaces, and 3. debugging, hyperparameter tuning. Nevertheless, they are provably secure.
Despite their non-trivial engineering, MPC programs enjoy strong security guarantees and relative ease of deployment. Even small organizations can deploy MPC without any specialized hardware. Thus, the algorithms developed and shared in SecureKL readily enable zero-leakage dataset-to-dataset evaluations before sharing data.
Assumptions
Consider high stakes domains where disparate data may have additive benefits to the existing data. To make privacy boundaries tractable, we make the following assumptions:
  1. 1.
    Existing knowledge is not private. The hospitals are aware of each other having such data to begin with. The hospitals may know of the available underlying dataset size and format, which is assumed to be uniform across the hospitals in the setup to simulate unit-cost. Hospitals frequently know of each other’s resources, and the available ICU units are contentious, not kept secret.
  2. 2.
    Uniformity of $|\mathcal{D}_{i}|$. Though each hospital gets to price their data and set their own budget, for generality, the uniformity assumption allows us to use the number of additional data sources $n$ as the main ”budget proxy” across different strategies.
  3. 3.
    Legal risks of sharing any data are omnipresent in high stakes domains. The risks with sharing sensitive data in data-leaking strategies, which we coin as $\pi_{d}$ (demographic distance) and $\pi_{s}$ (small sample), are not made explicit, but assumed to be ”moderate” and ”high” respectively. This abstraction side-steps legal discussion, which would go beyond the scope of our paper.
  4. 4.
    No malice is assumed on any of the parties involved, as each hospital wants to authentically sell their data and set up a potential collaboration. This assumption becomes stronger when the number of parties grows or when the setup changes to potentially more competitive industries with less trust. We note our limitations in Section 6.3.
4 Methods
Strategy Sub-strategies Leakage
$\pi_{0}$ Blind (baseline) n/a zero
$\pi_{p}$ Private (SecureKL) n/a minimal
$\pi_{d}$ Demographic sex, gender, race moderate
$\pi_{s}$ Subset Sampling 1%, 10%, 100% high
Table 1: Concrete selection strategies $\pi$, differentiated by leakage risks. For a strategy $p$, a set of targets is chosen, i.e., $T\leftarrow\pi(\mathcal{D}_{o},n)$.
4.1 Defining Dataset Acquisition Strategies
Given the private and non-private data acquisition strategies illustrated in Figures 1 and 3, we associate privacy leakage as a primary cost of data partnerships. This section formalizes them by their leakage risks, summarized in Table 1.
A, high leakage, sharing raw data. $\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered (Centers for Medicare & Medicaid Services, 2025). Though leakage can be controlled through $k$, the data is inherently sensitive. The underlying distance metric follows the $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score introduced by Shen et al. (2024) , which we define in the following subsection.
B, moderate leakage, sharing summary statistics. $\pi_{d}(n)$ uses demographic metadata to guide data selection. This is implemented through ratio distance between source and target distributions, which may be considered aggregates therefore potentially not sensitive, such as when the underlying aggregation function $\phi$ is differentially private.
C, zero/minimal leakage, sharing no additional information besides what is assumed public, and what our method outputs, such as a score or a ranking. There are two methods: a. Blind selection baseline: $\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. Prior works suggests that when $n=1$, randomly selecting a source in hospital ICU may be risky and inefficient. b. Our method $\pi_{p}(n)$ selects data sources based on privacy-preserving measure for data combination, specifically Private $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
4.2 KL-based Methods, Without Privacy
In subset-sampling strategy $\pi_{s}(n,k)$, each of the candidate entities will leak a set of raw data. $\pi_{s}$ is implemented with KL-based measures similar to Shen et al. (2024). Recall $\mathrm{KL}(P||Q)$ is not symmetrical, meaning that it is not a “metric” that satisfies triangle inequality. Intuitively, this means the measure is directional: a hospital’s distribution $P_{o}$ may be “close” to the target distribution $P_{i}$, but not the other way around:
$\mathrm{KL}(P_{o}||P_{i})=\int_{x\in\mathcal{X}}\log\frac{P_{o}(\mathop{}\!\mathrm{d}x)}{P_{i}(\mathop{}\!\mathrm{d}x)}P_{i}(\mathop{}\!\mathrm{d}x)$ (Ideal Estimator)
Because we only have access to finite data $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$, approximations are needed. Typically, a learned model can capture distributional information, used to estimate continuous entropy. Thus the joint distribution of features and labels from both the source and target are included, with the goal of deriving an efficient estimator for $\mathrm{KL}(P_{o}||P_{i})$ that captures distributional shift from source to target.
Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model (Cox, 1958) on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covariates. The goal is to infer dataset membership using “proxy labels”, defined as follows:
$\mathcal{I}(x,y)=\begin{cases}1&\text{if }(x,y)\in\mathcal{D}_{o}\\ 0&\text{if }(x,y)\in\mathcal{D}_{i}\end{cases}$
A binary predictor is fit on this combined dataset, predicting $\mathcal{I}$ from $(\mathcal{X},\mathcal{Y})$ using logistic regression. The model’s output $p(x,y)$, also called the probability score, is Score(x,y)(x,y).
A score of $0.5$ or less means the datasets are not distinguishable, making the data potentially useful. Shen et al. (2024) established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task, similar to Miller et al. (2021). We note again that even though this model is trained on both parties’ data, the final algorithm that the hospital uses to train on combined data is not restricted.
Then, the resulting model’s probability score function $\text{Score}(\cdot)$: $\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$ (KL-XY)
We reproduce Shen et al. (2024) ’s results that $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ is predictive of downstream change in AUC in hospital setting (Table 4). Let the score function $g_{\mathrm{KL}}$ be the approximate of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$. The strategy selects the most likely hospital with the closest distance under the measure:
$\pi_{s}(n=1,k=K)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{KL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$ (KL-based Strategy, in plaintext)
When only a subset is available, this function is adjusted by swapping $\mathcal{D}_{i}$ for $\mathcal{D}_{i}^{\prime}\subseteq\mathcal{D}_{i}$ where $|\mathcal{D}_{i}^{\prime}|=k$. We denote the full dataset size as $K=|\mathcal{D}_{i}|$.
Though leakage can be controlled through $k$, yet the data is inherently sensitive. In ICU data, simulate that a default of $1\%$ is shared, so $k=3000\times 1\%=30$, though we run experiments with $k\in\{3,30,300,3000\}$.
4.3 SecureKL: Private KL-based Method
Figure 4: Our method SecureKL. Each side encrypts their data. Then, a model is privately trained on their joint data. Afterwards, their divergence is computed. Finally, only the final result of this dataset-to-dataset evaluation is revealed.
Using MPC, we extend on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to require no information sharing (besides what was already assumed public). Specifically we leverage private tensors and secure gradient descent in CrypTen (Knott et al., 2021) to implement private $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. As illustrated in Figure 4, the logistic regression as well as the scoring need to be implemented in private.
Denote the private encoding of $x$ as $[x]$.
$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}([x,y])).$ (Secure KL-XY)
Let the score function $g_{\mathrm{SKL}}$ be the secure approximation of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$. The strategy selects the most likely hospital with the closest distance under the measure:
$\pi_{p}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{SKL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$ (SecureKL Strategy, encrypted)
As shown in Figure 4, any KL-based measure $g_{\mathrm{SKL}}$ can be adapted to our setup. We mainly use $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ as the underlying measure. Its performance is detailed in Section 6.2. Additionally, even though our implementation measures distance of data between one source and one target party, the setup readily extends to accommodating multiple parties. Section 7.2 discusses potential deployment challenges.
Figure 5: SecureKL: Overall Correctness. Rank correlation between SecureKL output and ground truth AUC change, $\delta_{i}$, from acquiring 11 additional dataset for a given source hospital $H_{o}$. We propose selecting data partner ranked by our secure system under $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score to reliably increase AUC gains. ($|\mathbf{H}|=12$ hospitals; colored by source.)
4.4 Trivially “Private” Baseline: Blind Selection
$\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. This random strategy may evade selection biases and help gather diverse data. Yet, prior work  (Shen et al., 2024) suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
4.5 Alternative Data-leaking Baseline: Sharing Summary Statistics
A relaxation to sharing no sensitive data is to share metadata. While demographic traits are often causal and available, their exact cause in relation to the task is not a priori established (without a highly effective model), therefore their success in distributional-matching is not guaranteed to be strong. Additionally, in practice, the most effective model that results from data combination may or may not be causally-sound. Nevertheless, we posit alternative strategy $\pi_{d}(n)$ to find the demographically close candidates to guide data selection: Let $\phi:\mathcal{D}\to\mathbb{R}^{m}$ be an $m$-dimensional summary statistic of a demographic feature i.e. the racial distribution of patients. Then, we use the distributional distance between $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$, characterized by their $L_{2}$-distance through $\phi$:
$\pi_{d}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,L_{2}(\phi(\mathcal{D}_{o})||\phi(\mathcal{D}_{i})).$ (Demographic-based Strategy)
5 Experiments
5.1 Experimental Questions
  1. 1.
    Consistency: Does MPC degrade the original measure’s effectiveness? Since MPC implementations introduce approximations, we first validate $\mathrm{SKL}$’s correctness by examining the correlation of private scores (SecureKL) vs. plaintext scores ($\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, with full data access). Then, for the consistency of choosing encryption in the hospital domains, we examine plaintext and encrypted versions’ correlation with downstream ground truth rankings across hospitals.
  2. 2.
    Positivity: Does our method pick entities that reliably improve performance? If source dataset $D_{o}$ can only add data from $n$ more sources, does our measure lead to eventual AUC improvements? Specifically, in multi-dataset combination, we examine whether using $\mathrm{SKL}$ can improve the source hospital’s downstream outcome. When selecting a single (or a few) additional data source, how many hospitals improve with our method? Additionally, we compare our method with alternative, privacy-leaking strategies.
  3. 3.
    Error analysis: If our privacy-preserving method is not the dominant strategy against alternatives including limited data accessibility, why? As we know, small and uncertain improvements for downstream tasks underscore the inherent difficulty of evaluating data utility without seeing the full data. We are especially interested in analyzing (a) hospitals with low $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ correlations, and (b) hospitals lagging AUC improvements using the random strategy $\pi_{0}$ or limited-sample strategy $\pi_{s}$.
5.2 MPC implementation
$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ includes training logistic regression model in private. We implement measures based on dataset divergence by building a custom logistic regression model over encrypted data leveraging CrypTen. The model parameters and input are encoded as 16-bit MPCTensors, ensuring that all computations, including forward passes, sigmoid activations, and gradient descent updates, are performed in private.
Additional baseline details
We additionally run our experiments on plaintext methods used in Shen et al. (2024), including the $\mathrm{KL}_{\mathcal{X}}$ measure, which is similar to $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ without using each data source’s labels:
$\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}(x)).$ (KL-X)
To compare against $\mathrm{KL}_{\mathcal{X}}$ for baseline, we additionally implemented its encrypted version, $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$, as a $g_{\mathrm{SKL}}$ candidate.
$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}([x])).$ (Secure KL-X)
Optimizers
Because L-BFGS – the optimizer prior work (Shen et al., 2024) used in plaintext-only with Scikit-Learn (Pedregosa et al., 2011) – is not available as an encrypted version, our MPC experiments are facilitated with SGD optimizer. A fair comparison between the scores obtained through plaintext and encrypted settings necessitates re-implementing plaintext scores, Score($X$) and Score($X,Y$), using logistic regression with SGD in PyTorch (Paszke et al., 2019). The hyper-parameter tuning for SGD in private and plain text are performed independently, as they do not transfer. Appendix 12 will discuss hyperparameter-tuning in detail.
5.3 Data and Model Setup
eICU dataset. The downstream task is $24$-hour mortality prediction from ICU data using the eICU Collaborative Research Dataset (Pollard et al., 2018). This dataset contains >200,000>200,000 real-world admission records from $208$ hospitals across the United States.
A note on data filtering. Because medical research is inherently complex, ensuring reproducibility on statistical methods can be significantly challenging.  Water et al. (2023) proposes that the research community follow a shared set of tasks with fixed preprocessing pipelines that are clinically informed – essentially a machine learning training and evaluation protocol on eICU – in order to facilitate method verification on the same benchmark. Therefore, our work follows their data cleaning criteria and the evaluation protocol. Additionally, we use the hospital exclusion criteria in (Shen et al., 2024) to obtain top $12$ hospitals, where the most patient visits are collected (each with >2000>2000 patients).
Downstream model and baselines for eICU. Recall that each strategy uses the same $K$ number of records per hospital – in our experiment, $K=3000$. For $\pi_{s}$ which leaks a subset $k$ of all samples, a default $k=1\%|\mathcal{D}_{i}|$ randomly drawn samples are shared. In ICU data, we run experiments on $\{0.1\%,1\%,10\%,100\%\}$.
Following holistic benchmarking tools in  (Water et al., 2023), our strategy comparisons take $1500$ samples and the downstream model performance – $\mathrm{AUC}_{o}$, $\mathrm{AUC}_{T}$ – uses $400$ samples (unless otherwise noted). Specifically, the AUC change, $\delta_{i}$ or $\delta_{T}$, comes from 1. combining $1500$ random samples from each selected dataset and 2. combining it with $1500$ samples from $\mathcal{D}_{o}$, and 3. subtracting the baseline model’s AUC [^2] The samples are fixed across all experiments, the sample numbers are chosen to match (Shen et al., 2024)’s setup..
Folktables dataset Though we primarily focus on hospital domain, we additionally validate using Folktables (Ding et al., 2021), predicting across 15 states an individual’s annual income exceeds \$50,000333Because the Folktables dataset contains simpler features and is low-dimensional, KL divergence is directly computed rather than estimated. For implementation details, see Appendix 15.. The details of our processing, which diverges from that of eICU, is included in Appendix 15.
6 Results and Analysis
We organize our results around three research questions: whether using multiparty implementation sacrifices $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$’s efficacy (consistency; Section 6.1), whether our method reliably picks hospitals that improve performance (positivity; Section 6.2), and where our method may fail (error analysis and limitations; Section 6.3).
6.1 Consistency Between Plaintext and Encrypted Computations
Because our encrypted computation is the first implementation of dataset divergence in MPC, we ought to show that $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ lead to highly comparable behaviors as $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}}$.
Spearman’s Rank Correlation Coefficient for Underlying Scores
For each source hospital $H_{o}$, use all full samples for $\mathcal{D}_{i}$. Between $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ on $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ for all remaining hospitals $H_{i}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.908$ with a range of $[0.691,1.0]$, obtaining $p<0.02$ across all hospitals. Between $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.9303$ with a range of $[0.455,0.991]$, with 11 of 12 hospitals achieving p-values below $0.05$. After applying Hochberg false discovery rate correction (Benjamini and Hochberg, 1995), our p-values remain significant. This range may be an artifact of sweeping hyperparameters independently in plaintext and encrypted optimisations, because sharing the same SGD hyperparameters would result in a tighter range. For all 12 hospitals, see appended Appendix 13 for details.
Impact of Adding Security on AUC Correlation
We further examine the effect by adding encryption through its impact on the downstream AUC, using how AUC improvements are ranked. This rank is compared with how secure measures (i.e., $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$) and plaintext measures (i.e. $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$) rank hospitals. This comparison investigates the extent of the shift in the full hospital ranking, when we switch from a plaintext setup to encrypted. For $H_{o}\sim\mathbf{H}$, we measure $\delta_{i}$ that results from adding $\mathcal{D}_{i}$ to $\mathcal{D}_{o}$ for all $i$. This correlates all target hospitals $\{H_{i}\}$ with their ground truths $\{\delta_{i}\}$ in the case of picking a single target hospital. We find the linear coefficient for encrypted $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be 0.182-0.182 and plaintext $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be 0.184-0.184 ($99\%$ matching). Both $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$ have a linear coefficient of 0.164-0.164 with $\delta_{i}$. For all strategies’ correlations with ground truth at $n=1$, see Appendix 11.
6.2 Positivity in Realistic Setup
We apply SecureKL in a pragmatic multi-source data combination problem, where each strategy acquires $n$ datasets for $n\in\{1,2,3\}$.
Overall Positivity.
For $n=1$, we find that $\pi_{p}$ improves AUC in 10 out of the 12 hospitals. When $n=2$ and $n=3$, we find that using $\pi_{p}$ consistently improves AUC for all hospitals. Overall, 34 out of the 36 dataset combinations we evaluate on have an AUC improvement $\delta_{T}>0$, suggesting that $\pi_{p}$ is a reasonable strategy for selecting hospital dataset combinations with a high expected return $\mathbb{E}[P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)]$ for the source hospital from using our strategy.
Demographic Summary Strategies Subset-sampling Strategies Secure Strategy (Ours)
Dataset $n$ $\pi_{d}$-gender $\pi_{d}$-age $\pi_{d}$-race $\pi_{s}(k=300)$ $\pi_{s}(k=30)$ $\pi_{p}\ \mathrm{Secure}\mathrm{KL}_{\mathcal{XY}}$
eICU 1 $\textbf{0.020}\pm 0.023$ $012\pm 0.016 0.0$$14\pm 0.015 0.0$$12\pm 0.014 0.0$$\pm 0.017 0.011±0.0$$1\pm 0.018$
2 $016\pm 0.016 0.0$$17\pm 0.017 0.0$$15\pm 0.013 0.0$$24\pm 0.020 0.0$$17\pm 0.019 0.0$$2\textbf{0.027}\pm$ 0.022
[…]
2 $\textbf{0.015}\pm 0.019$ $\textbf{0.015}\pm 0.014$ $0\pm 0.013 0.011±0.0$$1\pm 0.014$ $014\pm 0.018 0.0$$010\pm 0.017$
[…]
Table 2: AUC improvements in mean and standard deviation, across all source regions for each strategy $\pi$, for eICU and Folktables setups. $n$ denotes the number of candidate datasets added to the source dataset. The small gains and high variance from adding selected datasets highlight the precarious nature of assessing data value in the real world. Bold indicates highest AUC improvement per $n$. Note: Only $\pi_{p}\ \mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ is private.
Figure 6: AUC change $\delta_{T}$ over all strategies in eICU prediction (higher is better). Our private dataset evaluation strategy $\pi_{p}$ outperforms demographic-based strategy $\pi_{d}$ (left), and subset-sampling strategy $\pi_{s}$ for $k=300$ (10%) and $k=30$ (1%) (right), after combining source data with the top 33 candidates.
Figure 7: AUC change $\delta_{T}$ over all strategies in Folktables dataset prediction (higher is better). All strategies exhibit comparable distributions, after combining data from top 3 candidates. In a noisy domain, our method is stable: it neither excels nor penalizes against non-private strategies.
Comparing With Alternative Strategies
Other strategies – $\pi_{0}$, $\pi_{d}$, and $\pi_{s}$ – can also arrive at “positive” datasets. Comparing private method to other strategies at $n=3$, i.e, $\pi_{p}(n=3)$, we describe our results in Figure 7:
  1. 1.
    $\pi_{p}$ (our method based on $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$) has a median $\delta_{T}$ of $0.020$, and a standard deviation of $0.015$. Our results indicate that for 5050% of the hospitals, $\pi_{p}$ gives a $\delta_{T}>=.02$. Compared to other strategies, $\pi_{p}$ has the highest median, the lowest standard deviation, and it is one of two strategies that improves performance for all hospitals.
  2. 2.
    Demographic-based strategies underperform compared to $\pi_{p}$ on average. However, we observe that $\pi_{d}$-gender can be highly effective for a subset of hospitals, as it achieves the highest 75th percentile (Q3) of $0.033$ among all strategies. This indicates that for 25% of hospitals, $\delta_{T}\geq 0.033$. Despite this, $\pi_{d}$-gender has a lower median value of $0.012$ compared to $\pi_{p}$, exhibits a high standard deviation ($0.022$), and degrades the performance for certain hospitals. Similarly, $\pi_{d}$-age has a median of $0.014$, and $\pi_{d}$-race has a median of $0.008$, both lower than $\pi_{p}$’s median.
  3. 3.
    Plaintext small-sample strategies, $\pi_{s}$, outperform all demographic-based methods but slightly underperform relative to $\pi_{p}$. For instance, $\pi_{s}(k=300)$ has a median $\delta_{T}$ of $0.0178$, and although it achieves $\delta_{T}>0$ across all hospitals, it performs worse on average compared to $\pi_{p}$ and exhibits a higher standard deviation ($0.017$). $\pi_{s}(k=30)$ has a median $\delta_{T}$ of $0.0165$. Compared to other strategies, it has the largest standard deviation ($0.024$), and it degrades the performance for some hospitals.
In summary, our method $\pi_{p}$ achieves the highest AUC improvement on average with the lowest standard deviation, demonstrating more consistent improvements for all hospitals. While the average improvement of $\pi_{p}$ is small, demographic-based and plaintext small-sample strategies exhibit greater variability, with some strategies improving performance for specific subsets of hospitals but underperforming or degrading results in others.
6.3 SecureKL Error Analysis
Underlying Data Limitations
In high-stakes domains, data partnerships are expensive, but potentially detrimental – this forms a challenging landscape for evaluating methods on real-world data. Indeed, as shown in Table 2, the AUC gain is small across all strategies, and the variance is high. This suggests that 3 hospitals’ data is likely still too small for the general task to the robust explains limited AUC gains, highlighting the need to maximize samples for informative decisions. The key distinction, however, is that privacy-leaking methods (demographic, small sample) and blind baseline risk performance declines in many hospitals while SKL consistently improves downstream tasks more reliably than alternatives, across all hospitals, over multiple runs.
Figure 8: The correlation between AUC change $\delta_{i}$ and underlying score $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ affects $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$’s efficacy. In Hospital 420, the $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ scores identify beneficial data ($\rho<0$). For Hospital 243, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ anti-correlates ($\rho>0$).
Underlying Score Limitations
Data addition algorithms underpin the effectiveness of our method. Even if $\mathcal{D}_{o}$ obtains access to all the plaintext data, there is no guarantee that $\pi_{p}$ can correctly predict whether the data is useful. As seen in Figure 8, Hospital 243’s utility when acquiring another data set is badly correlated with plaintext and encrypted KL-XY scores. This leads to its bad strategy for acquiring the top 3 hospitals, as seen in the middle pane of Figure 9. Interestingly, for this hospital, no other informational strategy excels, either, so choosing a random 3 may be preferred.
This behavior stems from the underlying measure, not from adding secure computation: in Figure 8 and Figure 9, the encrypted performance closely follows that of plaintext performance, for both good and bad downstream correlations.
Sometimes, Not All Underlying Data Is Needed
Relatedly, when seeing a few samples can successfully identify useful candidate hospitals, $\pi_{s}$ (which is on small samples) outperforms $\pi_{p}$ (which is on full samples).
In the right panel of Figure 9, hospital 199, the smaller sample sizes achieve a score that better reflects ground truth as a data addition strategy. In that case, the hospital may not need the full sample to know which target hospitals to collaborate with.
This behavior is specific to the interaction of the data and the underlying score, and does not affect the general insight that adding private computation preserves privacy (and eases privacy-related risks that hinder data sharing). We further note that our method still clearly applies to encrypted computation on a smaller data set under data minimization.
7 Discussion
7.1 SecureKL Contribution
After establishing that $\pi_{p}$ with $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$is a robust strategy in practical downstream performance, we hereby summarize the benefits of $\mathrm{SecureKL}$ and elaborate on their practical implications.
Matching Plaintext Performance in Downstream Tasks. Our major contribution is to match plaintext performance with no data sharing. Using MPC provides input privacy, meaning that if both hospitals only want to know the resulting score, the computation can be done without leaking original data. This strong guarantee can significantly ease the tension related to privacy and compliance in setting up a collaboration, leading to a practical ”data appraisal stage” in data-limited high stakes domains.
Gain from Data Availability. In contrast to limited-sample approaches, a key advantage for our method $\pi_{p}$ is that it takes advantage of all the underlying data – generally impossible with non-secure methods for private data in heavily regulated domains. The general intuition is that data is localized; therefore, once a good target hospital is identified, we would prefer to acquire all of the data. It may be tempting to assert that we prefer the highest $k$ for data addition algorithms as well. In our experiments, while this is generally true, the smaller $k$ sometimes outperform larger $k$ in plaintext strategy $\pi_{s}$, which we investigate in Section 6.3 and in Figure 9. This occasionally non-monotonic behavior mirrors the challenge of data combination itself: even within one source dataset for the same estimator, more data is not necessarily better. This suggests the potential for a hospital-specific alternative to sharing a large amount of data for some source hospital, and points to future directions to using secure computation on a minimal-sized sample dataset for minimal performance overhead while remaining private.
Potential Improvements to SecureKL In the case where that output can be sensitive, i.e., when a source hospital queries a target hospital multiple times and accrues information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response (Dwork et al., 2014).
7.2 Potential Challenges to Broader Adoption
Our code is readily usable by small organizations. While our approach generalizes to model-based measures (by substituting $g_{\mathrm{SKL}}$) and scales to multiple parties, our work also uncovered deployment limitations.
  1. 1.
    Operational: engineering personnel limitations. While our implementation requires little cryptographic knowledge to deploy, it still needs technically-trained staff at each participating hospital to collaborate and maintain. This skill is similar to using pre-packaged software, cleaning data, and setting up network calls.
  2. 2.
    Engineering Extensions: Extending any MPC protocol is non-trivial, as security engineering is a specialized skill. While $\mathrm{SecureKL}$ applies broadly to other underlying scores in multi-party setups, validating a new MPC algorithm requires software engineering – prototyping, tuning, debugging – and numerical verification – akin to data analytics and research - likely requiring technical talents who can be especially costly for hospitals to retain in-house.
  3. 3.
    Framework Limitation: While CrypTen is designed to accommodate PyTorch, it is a research tool where not all plain text functionalities are implemented. For example, writing optimizers – such as L-BFGS – and custom operators that are not readily available requires both machine learning and cryptography knowledge. Moreover, the protocol incurs additional computational overhead, especially if hyper-parameters become more complex to sweep [^4] For our work, the performance metrics are provided in Appendix 14 for reference. This will likely improve with time, as new frameworks address design shortcomings.
  4. 4.
    Inherent to Secure Computation: When the method requires significant hyper-parameter tuning, such as using SGD on small batch data with learning rate schedules, plaintext tuning may not transfer perfectly. As detailed in Appendix 12, our hyperparameters for SGD differ in encrypted and plaintext settings. Thus, as encrypted computation hides loss curves and training details by default, development is expected to be complex. This is because both hospitals want to ensure model fit with secure evaluation, but may not want to expend the computational cost of private hyperparameter sweeping.
Figure 9: Mean AUC ($\pm$ standard deviation). Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{s}(k=30)$ and $\pi_{0}$. Middle: All strategies perform similarly. Right: $\pi_{s}(k=300)$ outperforms $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
8 Related Work
We briefly relate alternative approaches towards dataset privacy for sharing. More work on data pricing, differential privacy, and private fine-tuning are included in Appendix 16.2.
Augmenting existing data with synthetic data in medical domains
Synthetic data generation has emerged as a promising approach to expand training datasets while preserving privacy. Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data (Goncalves et al., 2020), medical imaging data (Thambawita et al., 2022), and electronic health records (Baowaly et al., 2019). These methods preserve statistical properties of the original data while providing differential privacy guarantees. Transforming data into a similar form that desensitizes certain attributes can be desirable (Drechsler, 2011; Howe et al., 2017; Nikolenko, 2021; Gonzales et al., 2023; Sweeney, 2002). Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself (Jordon et al., 2021). Additionally, outside the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks (Narayanan and Shmatikov, 2006; Jordon et al., 2021).
In addition, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables (Goncalves et al., 2020). For tabular data, methods like CTGAN and TVAE (Xu et al., 2019) have demonstrated ability to learn complex distributions while preserving correlations between features. However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance (Assefa et al., 2020). Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees (Jordon et al., 2018). While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset (Yang et al., 2024a).
Secure Data Combination
Recent work has explored methods for securely combining datasets while preserving privacy and improving model performance. Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data (Aono et al., 2017). However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets (McMahan et al., 2017). More recent techniques have introduced frameworks for evaluating potential data partnerships before commitment. These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets (Leung et al., 2019; Chakraborty et al., 2024). Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information (Duan et al., 2021). Others trained the downstream model in private, but limited to LASSO (van Egmond et al., 2021). Several systems have been developed to facilitate secure data combination in specific domains. In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance (Raisaro et al., 2018; van Egmond et al., 2021). Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality (Liu et al., 2021).
Federated Learning
Cross-silo federated, decentralized, and collaborative ML (McMahan et al., 2017; Li et al., 2020; Bonawitz et al., 2019; Kairouz et al., 2021) focus on acquiring more data through improved data governance and efficient system design. Healthcare machine learning is considered especially suitable, as health records are often isolated (Rieke et al., 2020; Xu et al., 2021; Nguyen et al., 2022; Cho et al., 2025). Yet, even though no raw data is shared, model parameters or gradients flow through the system. As the federated computing paradigm offers no privacy guarantee, the system is vulnerable to model inversion (Geiping et al., 2020) and gradients leakage attacks (Boenisch et al., 2023; Zhu et al., 2019). A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimization is traded off with privacy (Lyu et al., 2022; Raynal and Troncoso, 2024). Building on the insight that useful data is often disparately-owned, we tackle the specific incentive problem between pairs of players where one side trains the model, instead of scaling up the number of parties through a federation.
Compared to vanilla Federated Learning, an MPC system (Shamir, 1979; Yao, 1982; Bonawitz et al., 2017; Knott et al., 2021) provides stronger guarantee in terms of input security. Model owners and data owners can potentially federate their proprietary data, including model weights, training, and testing data, can work together under stringent privacy requirements. Our work extends the line of work by  (Xu et al., 2022; Yang et al., 2024b; Bonawitz et al., 2017) that demonstrates the potential of incorporating MPC in various federated scenarios. On the practical side, unlike mobile-based networks for secure federated learning protocols (Bonawitz et al., 2019), our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
Comparing with Federated Learning with Privacy Guarantees
Privacy incentivizes federation (Usynin et al., 2024). Specifically, preserving privacy between the parties under federated learning uses secure computation methods (Truex et al., 2019; Bonawitz et al., 2017; Stripelis et al., 2021; Cho et al., 2025; Froelicher et al., 2021) combined with differential privacy (Xu et al., 2023; Usynin et al., 2024; Avent et al., 2017), trusted execution environments (Mo et al., 2021) – an approach that is coined ”Privacy-in-Depth” in Kairouz et al. (2021).
Performing dataset evaluation with MPC, as in SecureKL, can be seen as an extension of data federation, by adding a separate privacy-preserving component before all parties commit to an entire federated learning system. It is low-commitment, as privacy is preserved by default, and there is no requirement to continue in any system; it tackles the incentives’ problem blocking collaboration. By demonstrating utility, the two parties foster trust. Our work complements the line of ambitious systems that incorporate MPC in potential federated scenarios where the participants are semi-honest (Yang et al., 2024b; Bonawitz et al., 2017; Cho et al., 2025).
9 Conclusion
Our work demonstrates that privacy-preserving data valuation can help organizations identify beneficial data partnerships while maintaining data sovereignty. Through SecureKL, we show that entities can make informed decisions about data sharing without compromising privacy or requiring complete dataset access. As the AI community continues to grapple with data access challenges, particularly in regulated domains like healthcare, methods that balance privacy and utility will become increasingly critical for responsible advancement of the field. As noted in Section 6.3, our approach has several limitations, including the fact that, despite impressive aggregate results, our method is less effective for individual hospitals, which motivates future work. Our method assumes static datasets and may not generalize well to scenarios where data distributions evolve rapidly over time. A sequential version of our framework may more closely model dynamic data collaborations. Future work should explore extending these techniques to handle more complex data types and dynamic distribution shifts while maintaining strong privacy guarantees.